OpenStudy (anonymous):

Find the sum of ((-ln 3)^n) / n!

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} (\frac{ (-\ln 3)^n) }{ n! } \]

OpenStudy (anonymous):

is not (inside the sum notation) \[\frac{ 1 }{ (\ln3)^n*n! }\]?

OpenStudy (anonymous):

this converges, clearly using ratio test or integration. But how do we get the sum??

OpenStudy (anonymous):

I don't know!! don't answer, let me have a chance to review my knowledge, I studied. I go now, kid